# references.bib

### v2

@@ IV-C Ablation Studies
TABLE III: Ablation studies for Image-Event fusion models
  Knowledge Transformer Feature Fusion Module Trainable Paras (M) Baseline-1 None Cross-Attention 49.2 Baseline-2 None FreDFuse 51.6 Baseline-3 One-stage FreDFuse 6+2.7 FUSE PST FreDFuse 1.4+4.7+2.7
We perform ablation experiments using the ViT-Small model as the backbone on the MVSEC dataset, as outlined in Tab.III. In Baseline-1 and Baseline-2, the model is trained from scratch on image-event pairs with depth ground truth. Baseline-1 uses two cross-attention mechanisms for feature fusion, while Baseline-2 incorporates the proposed FreDFuse (Sec. III-D). Baseline-3 combines FreDFuse with a one-stage knowledge transfer approach, initializing the event encoder with image-based foundation model weights and training both the event encoder and FreDFuse simultaneously.
Tab. IV and Tab. V present the results of the ablation experiments on the MVSEC dataset, specifically in the outdoor_day1 and outdoor_night1 scenes. The results demonstrate that both components of our proposed method, PST and FreDFuse, positively influence model performance.
@@
V Conclusion
In this paper, we propose FUSE, a novel framework that addresses two fundamental challenges in image-event joint depth estimation: cross-modal knowledge transfer under extreme data scarcity and physics-driven frequency conflicts during feature fusion. By aligning image-event latent spaces through a two-stage adapter tuning strategy, PST successfully transfers geometric priors from image depth foundation models to the joint modality domain without requiring any depth-labeled image-event pairs. This proves that event-depth representations can be effectively bootstrapped from image counterparts when guided by physics-aware constraints, opening new possibilities for data-efficient cross-modal learning. Unlike conventional fusion heuristics, the proposed FreDFuse explicitly resolves the spectral incompatibility between static images (low-frequency dominance) and dynamic events (high-frequency bias) through Gaussian-Laplacian pyramid decomposition. The FUSE framework achieves SOTA performance on both real-world (MVSEC) and synthetic (DENSE) benchmarks while enabling lightweight deployment.
While FUSE demonstrates strong zero-shot generalization, two limitations warrant further investigation. Current fusion treats events as fixed-interval histograms, potentially underutilizing their microsecond temporal resolution. The framework assumes synchronized image-event pairs, which may not hold for all hardware setups. Future work will explore event-guided dynamic fusion that adaptively weights frequency branches based on motion intensity, and unsupervised cross-sensor calibration to relax synchronization requirements. These extensions could further solidify image-event depth estimation as a universal solution for extreme environment perception.
TABLE IV: Quantitative results of the ablation experiments on the MVSEC[mvsec] outdoor_day1
@@ V Conclusion
TABLE V: Quantitative results of the ablation experiments on the MVSEC[mvsec] outdoor_night1
@@
[…]

### v3

@@ IV-C Ablation Studies
TABLE IV: Ablation studies for Image-Event fusion models
  Knowledge Transformer Fusion Module Paras (M) Baseline-1 None Cross-Attention 49.2 Baseline-2 None FreDFuse 51.6 Baseline-3 One-stage FreDFuse 6+2.7 FUSE PST FreDFuse 1.4+4.7+2.7
We conduct ablation studies on the MVSEC dataset using the ViT-Small backbone (Tab. IV). Baseline-1 and Baseline-2 are trained from scratch with depth supervision; the former uses cross-attention for fusion, while the latter adopts our proposed FreDFuse (Sec. III-D). Baseline-3 adds one-stage knowledge transfer by initializing the event encoder with image-based foundation model weights and training it jointly with FreDFuse. Tab. V and Tab. VI report results on outdoor_day1 and outdoor_night1. Both PST and FreDFuse contribute significantly to performance gains.
@@
TABLE V: Quantitative results of the ablation experiments on the MVSEC [33] outdoor_day1
@@
TABLE VI: Quantitative results of the ablation experiments on the MVSEC [33] outdoor_night1
@@
IV-D Computational Overhead Analysis
We analyze the computational overhead of our FUSE framework in terms of parameters, GFLOPs, and inference latency, using a ViT-L backbone and $266\times 266$ input resolution on a single RTX 3090 Ti GPU. We compare three models: (1) an image-only baseline (Depth Anything V2), (2) a dual-stream baseline with standard cross-attention (Baseline-1), and (3) our FUSE with FreDFuse.
As shown in Table VII, FUSE roughly doubles the parameters and GFLOPs compared to the image-only baseline due to its dual-encoder design, a common trait in multimodal approaches. Compared to Attention Fusion, FreDFuse introduces only a small overhead: +2.56% parameters, +0.46% GFLOPs, but a higher latency (+24.26%), mainly due to its pyramid-based decomposition, which adds memory and synchronization overhead despite low FLOP cost.
V Conclusion
We present FUSE, a framework for image-event joint depth estimation that tackles two challenges: cross-modal knowledge transfer and spectral mismatch during fusion. A two-stage adapter tuning transfers priors from image depth models to the joint domain without labeled image-event pairs. The proposed FreDFuse module leverages Gaussian-Laplacian pyramids to align frequency components from the two modalities. FUSE achieves state-of-the-art performance on MVSEC and DENSE benchmarks.
However, FUSE has several limitations. First, converting asynchronous events into fixed-interval frames reduces temporal fidelity, missing fine motion details. Second, it processes samples independently, ignoring inter-frame dynamics and temporal consistency. Finally, its relatively high latency—due to dual-stream processing and multiscale fusion—limits real-time deployment. Future work should focus on two main aspects. The first is to explore native asynchronous processing and recurrent architectures to better exploit temporal dynamics and improve depth consistency. The second is to optimize the fusion module and investigate model compression techniques, which are crucial for reducing latency and enabling real-time deployment.
TABLE VII: Efficiency Comparisons Between FUSE and Baselines.
  Model Paras (M) GFLOPs Latency (ms) Image-Only Baseline 335.32 309.61 29.20 Attention Fusion 660.05 549.21 57.90 FUSE 676.92 551.74 71.95
ACKNOWLEDGMENT
The research was supported by the National Natural Science Foundation of China (U23B2009, 62471158).